# Almost Surely Safe Alignment of Large Language Models at Inference-Time

### v2

@@ D.3 Critic Network and Training Process
Training Process
The critic network is trained using a combination of optimization techniques aimed at predicting the safety compliance and task cost accurately. The network is optimized with the hyperparameters in Table 3.
@@
D.4 Qualitative Comparisons
We present several examples from the PKU-SafeRLHF test dataset to demonstrate how InferenceGuard steers the base model outputs towards safer responses. These examples highlight differences in response safety between InferenceGuard and baseline methods, including BoN, Beam Search, RECONTROL, and ARGS, evaluated on the Alpaca and Beaver-v3 models, as shown in Figure 4 and 5. In each case, InferenceGuard successfully adheres to safety constraints while maintaining task performance.

### v3

@@ D.5 Critic Network and Training Process
Training Process
The critic network is trained using a combination of optimization techniques aimed at predicting safety compliance and task cost accurately. The network is optimized with the hyperparameters in Table 10.
@@
D.6 Qualitative Comparisons
We present several examples from the PKU-SafeRLHF test dataset to demonstrate how InferenceGuard steers the base model outputs towards safer responses. These examples highlight differences in response safety between InferenceGuard and baseline methods, including BoN, Beam Search, RE-Control, and ARGS, evaluated on the Alpaca and Beaver-v3 models, as shown in Figure 6 and 7. In each case, InferenceGuard successfully adheres to safety constraints while maintaining task performance.